Continuous Interest
10. If you invest $7,000 at 12 % compounded continuously, how much world it be worth in three
years?
Answer #1

An amount P invested at r % continuously compounded for t years becomes Pe ^(rt).

So, A = Pe^(rt) = 7000*e ^(0.12*3) = 10033.3059